find the domain of the composite function fog
f(x)= x^2+8; g(x)= sqrt.x-1

Answer by stanbon(75887)

find the domain of the composite function fog
f(x)= x^2+8; g(x)= sqrt.x-1
-----
fog = f[g(x)] = f[sqrt(x-1)] = (x-1)+8 = x+7
-----
Domain: All Real Numbers >= 1.
===============================
Cheers,
Stan H.